# Pose-Diversified Augmentation with Diffusion Model for Person Re-Identification

### v1

@@ 3.4 Conditional Diffusion Models for Pose-Diversified Augmentation
4 Experiments
4.1 Implementation Details
We utilize ResNet50 [18] as the feature backbone for our Re-ID model, initialized with weights pre-trained on ImageNet1K [44]. The training of our generative model is conducted in two stages: initially learning a general human representation using a fashion video dataset [60], followed by fine-tuning the model with person Re-ID datasets. Throughout both stages, the weights of the autoencoders [59] and CLIP [41] image encoder are frozen, focusing on the reference U-Net, denoising U-Net, and pose guider.
In the first stage, the reference U-Net and denoising U-Net are initialized from the original Stable Diffusion [42] model and fine-tuned with the fashion video dataset [60]. The fashion video dataset [60], comprising 192,402 frames from 500 human videos. All input images are resized to 768$\times$768 with a batch size of 2. Mean Squared Error (MSE) loss is calculated for each pair of a generated images, optimized using Adam [29] with a learning rate of 1e-5 and weight decay of 0.01. Training for 30,000 iterations takes 15 hours on one NVIDIA RTX A6000 GPU.
In the second stage, the generative model pre-trained on the fashion video dataset [60] is fine-tuned with Re-ID datasets consisting of pedestrian images from CCTV cameras. Images are resized to 192$\times$384, we increased the batch size to 4. The training process is the same as in the previous stage. It takes about 6-7 hours for the DukeMTMC-reID [71] dataset with 16,522 training images and the Market1501 [65] dataset with 12,936 training images on one NVIDIA RTX A6000 GPU.
Our work employs data augmentation via a generative model to enrich the training data for the Re-ID model. We start by randomly sampling 20 target pose guidances from the Everybody Dance Now dataset [4], an out-of-domain dataset not used in the training of the Re-ID model. Next, we randomly sample between 5 to 7 instances from each frame sequence of person identity (pid) in Re-ID datasets and input these instances into the generative model. The generative model then produces 65 to 70 images for each sampled instance, resulting in an generated dataset. From this generated dataset, which consists of 80,000 images, we randomly sample approximately 45,500 and 49,000 images for Market-1501 and DukeMTMC-reID, respectively. These generated images are then added to the original training datasets to create the augmented training sets. For the DukeMTMC-reID [71] dataset, which originally consists of 16,522 training images from 702 pids, the addition of approximately 49,000 generated images results in a new training set comprising 30,522 images from 702 pids. Similarly, for the Market-1501 [65] dataset, which originally consists of 12,936 training images from 750 pids, the addition of approximately 45,500 generated images results in a new training set comprising 26,936 images from 750 pids. This augmentation strategy leverages the generative model to significantly increase the diversity and quantity of the training data, thereby enhancing the model’s ability to generalize and perform effectively on the Person Re-identification task.
4.2 Experimental Details
We perform experiments on two publicly available person Re-ID datasets, i.e., Market-1501 [65] and DukeMTMC-reID [71]. We present results using two standard Re-ID metrics: cumulative matching characteristics at Rank-1 (R1) and mean average precision (mAP).
We trained our Re-ID model with augmented Re-ID datasets: Market-1501 [65] and DukeMTMC-reID [71]. For each training iteration, the batch size of pids is set to 32 and the number of instances per pid is set to 8, resulting in a total batch size of 256. We utilize Triplet loss [22] and classification loss to distinguish different person identities. The triplet loss [22] is calculated by first obtaining the euclidean distance between each instance using the 2,048-dimensional feature extracted from ResNet50 [18]. The objective is to ensure that the hard-negative distance, the shortest distance between instances of different person identities, is larger than the hard-positive distance, the longest distance between instances of the same person identity. The margin of the triplet loss [22] is set to 0.5, meaning the gap between the hard-negative distance and the hard-positive distance should be larger than 0.5. In addition to the triplet loss [22], we use a classification loss, which is a CrossEntropy [62] loss computed between the pid and the output logits from the classifier. The total loss for training is the sum of the triplet loss and the classification loss. We apply the Adam [29] optimizer with different learning rates for ResNet50 [18] and the classifier. For the ResNet50 [18] backbone, we set the learning rate to 1e-5, and for the classifier, we set the learning rate to 1e-4. The training process with 20,000 iterations takes approximately 2 hours on two NVIDIA RTX A6000 GPUs.
Table 1: Quantitative validation of our design choices.
Diverse Human Pose Diverse Viewpoint Pre-trained Market-1501 DukeMTMC
Augmentation Augmentation Diffusion mAP R1 R5 mAP R1 R5
(I) 68.5 84.9 92.4 55.5 73.6 85.4
(II) 68.6 85.0 92.5 55.8 73.4 85.2
(III) 69.5 86.2 93.9 60.5 77.4 87.2
(IV) 69.4 86.1 93.6 60.1 77.2 86.9
(V) 67.6 84.5 93.4 54.1 72.2 83.8
(VI) 72.6 87.4 94.5 63.2 78.7 88.5
4.3 Ablation Study and Analysis
In Table 1, we conduct an ablation study to verify our design choices: 1) human pose augmentation, 2) camera viewpoint augmentation, and 3) the effectiveness of the pre-trained diffusion model.
(I) is a baseline Re-ID model trained on the original dataset, where Diff-ID is not used for the augmentation. For (II), we train a diffusion model on the training dataset and augment the data using random human poses and viewpoints within the training distribution. The marginal performance gain from (I) to (II) demonstrates that augmentation within the training distribution is insufficient. For (III) and (IV), we augment the training dataset using diverse human poses from an external dataset and uniformly sampled viewpoints, respectively. Both augmentations show significant performance gains. Notably, for the DukeMTMC-reID dataset [71], which exhibits a stronger bias in camera viewpoints, we observe improvements of 4.7%p and 4.3%p for each augmentation strategy. This validates the effectiveness of targeting the biased distributions of human pose and viewpoint for augmentation. In (V), we use the same architecture as Diff-ID, except for the initialization, where we randomly initialize the SD model instead of using its pre-trained weights. We found that with random initialization, the performance is even worse than the baseline (I). We believe this is because a diffusion model trained from scratch has limited ability to generalize to poses or viewpoints rarely seen in the training set. This result supports our decision to leverage the pre-trained knowledge of SD. (VI) demonstrates the final version of our model. The performance gains from both (III) and (IV) prove that both types of augmentation are useful and have a complementary effect, leading to further performance improvements.
Table 2: Quantitative comparison of different settings. We gradually increase the ratio of generated images and compare the results of various settings. Each row represents the result for an approximate number of augmented images.
# of Images Market-1501 DukeMTMC
mAP R1 mAP R1
3,500 69.8 86.4 57.3 75.0
7,000 71.1 86.3 60.0 76.5
10,500 71.6 87.0 61.2 77.2
14,000 71.1 86.4 61.2 77.6
17,500 71.5 86.2 62.1 77.2
21,000 71.5 86.0 62.5 78.3
24,500 72.1 86.6 62.5 77.5
28,000 72.0 86.7 62.2 77.8
31,500 72.2 87.0 62.3 77.7
35,000 72.2 87.2 62.1 77.8
38,500 72.2 86.8 62.4 77.3
42,000 72.1 86.9 62.6 78.1
45,500 72.6 87.4 62.8 77.6
49,000 72.4 86.6 63.2 78.7
52,500 72.4 87.0 62.9 78.1
56,000 72.4 86.7 62.4 77.6
In Table 2, we present an additional ablation study to determine the optimal number of images used in augmentation. The proposed settings aim to evaluate the effect of increasing the training dataset size. We randomly sample between 5 to 7 instances from each frame sequence of person identity (pid) in the Re-ID datasets and input these instances into the generative model. The generative model produces a minimum of 5 to a maximum of 80 images for each sampled instance. We then incorporate varying numbers of these generated images to the original dataset.
We expected to see higher results as the size of the augmented dataset increased. However, in both cases using different datasets, namely the DukeMTMC-reID dataset [71] and Market-1501 [65], there is an optimal number of images that enhance performance. For Market-1501 dataset, the highest result is obtained when approximately 45,500 images are added. For DukeMTMC-reID dataset, the highest result is obtained when approximately 49,000 images are added. This result demonstrates the high quality of the generated images, as adding a significant number of images still results in high performance.
Table 3: Quantitative comparison on standard Re-ID benchmarks. Note that the Re-ID Experts in the first row group are not directly comparable, as our primary focus is on dataset generation. For augmentation-based methods, we train the same Re-ID model on the datasets generated by each method to ensure a fair comparison. *: The authors did not provide a pre-trained model.
Methods Market-1501 DukeMTMC
mAP R1 mAP R1
Re-ID Experts
LTReID [52] 89.0 95.9 80.4 90.5
DRL-Net [25] 86.9 94.7 76.6 88.1
BPB-ReID [47] 89.4 95.7 84.2 92.4
AAformer [74] 87.7 95.4 80.0 90.1
CLIP-ReID [31] 90.5 95.4 83.1 90.8
Augmentation-based Methods
FD-GAN [12] 68.1 85.1 51.3 69.3
Xing-GAN [49] 70.3 86.1 - -
Diff-ID (Ours) 72.6 87.4 63.2 78.7
Figure 5: Qualitative comparison. We qualitatively compare the generated output to FD-GAN [11] and XingGAN [49]. Our method demonstrates significantly better fidelity while faithfully capturing the identity of the person in the reference image and accurately following the target pose.
4.4 Quantitative Comparisons
In this section, we evaluate the performance of our method, and compare it with the augmentation-based Re-ID methods. The results are shown in Table 3. Note that our augmentation method can be applied to arbitrary Re-ID models, and the performance can be improved when better Re-ID method is applied.
To ensure a fair assessment and maintain consistency, the same Re-ID model architecture was used to train on datasets produced by each augmentation method. Compared to previous augmentation methods, our augmentation method demonstrates significantly improved performance. This result also validates our augmentation framework, highlighting the effectiveness of filling in the sparsely distributed poses and viewpoints. Note that the first row group, Re-ID Experts, is not directly comparable to our methods, as our primary focus is on addressing biases in the training dataset rather than developing superior Re-ID models.
Figure 6: Visualization of generated data. Given a reference image, we sample five images generated by our method. These outputs demonstrate the model’s capability to produce diverse and realistic variations. Ref. denotes the reference image.
Figure 7: Generated results under comprehensive conditions. Using a reference image of individuals carrying backpacks, we generate back views using our method. This demonstrates that our method can be effectively applied under complex and potentially confusing conditions.
4.5 Qualitative Results
In Figure 5, we present a qualitative comparison to GAN-based augmentation methods [33, 69, 70]. Unlike GAN-based methods, Diff-ID effectively generates poses sourced from outside the dataset. GAN-based methods struggle to generalize to diverse poses, resulting in blurry outputs and a limited ability to main the identity of the reference and the target pose, especially when dealing with complex elements like accessories or specific clothing details. This issue is particularly pronounced when generating images of people carrying bags. GAN-based methods often fail to accurately distinguish between bag and the person’s clothing, leading to confusion. For example, they might generate person’s top in the color of the bag from the reference image, or alter the shape and color of the bag, or even change the entire outfit based on the bag’s color, as shown in Figure 5. This confusion is especially prevalent when the reference image features a backpack, which can cover the person’s entire back in the back view, making it easily mistaken for a top. This is due to GAN-based models’ lack of general knowledge about the world.
In contrast, large-scale generative models possess comprehensive knowledge about humans, their clothing, and accessories. They understand the context of outfits and generate images that preserve the identity of the reference image. This advantage underscores our decision to utilize a large-scale pre-trained diffusion model, which can generalize to poses and viewpoints underrepresented in the training dataset. Its extensive pre-training on a diverse range of data enables it to handle these complexities more effectively.
4.6 Visualization of Generated Data
In Figure 6, we present visualization results on two datasets, DukeMTMC-reID [71] and Market-1501 [65]. Given a reference image, our method faithfully preserves its identity while being able to generate diverse poses with high fidelity.
Given a reference image, our method leverages the strengths of a large-scale pre-trained diffusion model to maintain the intricate details of the subject’s identity. This includes preserving clothing details, and other distinguishing characteristics, ensuring that the generated images are not only visually coherent but also identity-consistent with the reference image.
In the DukeMTMC-reID [71] and Market-1501 [65] dataset, our method excels in generating various poses while preserving the unique identities of the individuals. Our method demonstrates its capability to generate high-fidelity images that faithfully reproduce the identities of the reference subjects across a range of poses and viewpoints.
Furthermore, our approach, facilitated by the large-scale pre-trained diffusion model, showcases an enhanced understanding of contextual details, such as the interaction between the person and their clothing or accessories by facilitate large-scale pre-trained diffusion model. This allows our model to generate more realistic and contextually accurate images. For instance, when the reference image includes a person carrying a bag, our model accurately distinguishes between the bag and the person’s clothing, ensuring that both elements are rendered correctly in the generated images. More qualitative results about this is shown in Figure 7.
For the additional qualitative results, please refer to Appendix A.
5 Conclusion
In this paper, we proposed Diff-ID, a novel data augmentation approach that leverages pre-trained diffusion models to diversify pose and viewpoint distributions in person re-identification training datasets. By sampling from outside pose sources and uniformly distributed camera viewpoints, we augmented Re-ID datasets with realistic images of individuals in rare poses and from uncommon camera angles. This augmentation strategy enables the training of Re-ID models that are more robust to variations in human pose and camera viewpoint. The key to Diff-ID’s success lies in its ability to integrate pose, viewpoint, and identity conditions into large-scale pre-trained diffusion models like Stable Diffusion [42], effectively leveraging the vast knowledge embedded in these models for generating high-quality augmented data. Comprehensive experiments on the Market-1501 [65] and DukeMTMC-reID [71] benchmarks demonstrated the effectiveness of our approach, with Diff-ID achieving significant performance improvements compared to baselines without data augmentation. Ablation studies further verified the importance of both diverse pose augmentation and camera viewpoint augmentation in enhancing the generalization capabilities of Re-ID models.
Appendix A Additional Qualitative Results
In this section, we present additional qualitative results produced by our framework, Diff-ID. Each subsection highlights different categories of generated images.
A.1 Generated images with of out-of-domain human pose distributions
Given a reference image in the training dataset, which is DukeMTMC-reID [71] and Market-1501 [65], we sampled various human poses from the Everybody Dance Now dataset [4]. We generated images that preserve the identity of the reference image while adopting human poses sourced from outside the training dataset. Fig. 8 demonstrates that the proposed model exhibits noticeable generalization ability and high-quality image generation.
Figure 8: Generated images with out-of-domain human poses. We show the generation results with diverse human poses.
A.2 Generated images with various camera elevations
We present the generation results with human poses captured at both high and low camera elevations. The human poses used as conditions for the generative model are extracted from the training dataset, with high elevation poses being relatively challenging. Fig. 9 illustrates how appearance and human poses vary with camera positions, particularly with changes in elevation, emphasizing the importance of data augmentation with diverse camera poses.
Figure 9: Generated images with various elevations. We generate images with high and low camera elevations. Ref. denotes the reference image.
A.3 Generated images with diverse poses
We generate images with diverse pose and identities for both Market-1501 [65] and DukeMTMC-reID [71] dataset. Fig. 10 demonstrates our framework’s capability to generate high-fidelity images that reproduce the identities of the reference images despite the diverse variations.
Figure 10: Generated images with diverse poses. We show additional qualitative results of the generated images. Ref. denotes the reference image.
A.4 Generated images with same poses
We generate images with same human poses and different identities for DukeMTMC-reID [71]. Diff-ID generates realistic images regardless of how human pose conditions are different from human poses of reference images. In Fig. 11, the first column of generated images demonstrates our framework’s capability to generate images despite extreme human poses.
Figure 11: Generated images with same poses. We generate images with same human pose and different identities. Ref. denotes the reference image.

### v2

@@ Conclusion
Additional qualitative results
Figure 8: Additional qualitative results. Examples of generated images from the Pose-dIVE augmented datasets. The results demonstrate realistic rendering while preserving the identity of the reference images and aligning accurately with the target poses.
In the main paper, we present generated images of Pose-dIVE augmented datasets using MSMT17 and Market-1501, showcasing the effectiveness of our approach. In this section, we provide additional qualitative examples from the MSMT17, Market-1501, CUHK03 (D), and CUHK03 (L) datasets.
The generated images in the Pose-dIVE augmented datasets exhibit several key strengths. Firstly, they maintain a high level of realism, effectively capturing the nuances of natural human appearances. Secondly, successfully preserves the identity of individuals from the input reference images, ensuring that the augmented data remains faithful to the original person’s characteristics. This is particularly important for re-identification tasks where identity preservation is critical. Thirdly, the generated images accurately follow the provided target poses, demonstrating the robustness in handling diverse and complex human poses. These additional results further validate the effectiveness of our proposed augmentation method across different datasets.
Analysis on the number of generated images.
Figure 9: Impact of the number of generated images per PID. Experiments are conducted in the Pose-dIVE augmented CUHK03 (L) dataset. We use CLIP-ReID baseline.
Figure 9 illustrates the effect of varying the number of generated images on the performance of the Re-ID model when using our proposed augmentation strategy. To evaluate this, we gradually increased the number of generated images in the training dataset and trained the CLIP-reID model, carefully observing how performance changed with each increment. It’s important to emphasize that this augmentation is only applied to the training dataset, with the test dataset remaining unchanged throughout the experiments.
The analysis reveals a clear trend: there is an optimal number of generated images that maximizes the performance of the Re-ID model. This trend was consistently observed across all the datasets we analyzed, allowing us to fix the ratio of generated to original images regardless of the specific dataset.
Visualizing Conditioning in Generative Models
Figure 10: Example SMPL, skeleton, depth and normal maps from external dataset. Examples of generated images from the Pose-dIVE augmented datasets. The results demonstrate realistic rendering while preserving the identity of the reference images and aligning accurately with the target poses.
We present visualizations of conditions used in our generative model, specifically, the skeleton, depth, and normal maps, in Figure 10. With these conditions, we can control the human pose and viewpoint of the images generated by our diffusion models.